OpenStudy (anonymous):

in a parking lot, 1 out of every 8 cars is blue,what percent of the cars in this parking lot are blue? The answer is 12.5% but can you explain how to do this.

OpenStudy (anonymous):

percentage of cars that is blue = \[\frac{1}{8} \times 100\]

OpenStudy (anonymous):

100/800

OpenStudy (anonymous):

\[\frac{100}{8}\rightarrow\] remember numberator with numerator denominator with denominator

OpenStudy (anonymous):

\[\frac{1}{8} \times \frac{100}{1}\]

OpenStudy (anonymous):

so i multiple 1*100 and 8*1which =100/8 then divide

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np:}

OpenStudy (anonymous):

you can also just divide 1 by 8, you get .125, make it a percent by moving decimal over, you get 12.5%